# Object Agnostic 3D Lifting in Space and Time

### v1

@@ 3.2.3 Space encoder
The space encoder uses the features from our motion encoder to model the relationships among joints in a single frame. We found the hybrid graph-based approach of [7] to perform favorably for our task. Given the motion features $\mathbf{F}_{M}\in\mathbb{R}^{T\times J\times D}$, a single space-layer has two simultaneous processing streams, one for capturing the local connectivity between joints $G_{\text{local}}$, and another for capturing global connectivity $G_{\text{global}}$. These two streams are concatenated and projected to provide an output containing a combination of both streams $\mathbf{F}_{S}$ as:
@@
The local graph-attention $G_{\text{local}}$ utilises an adjacency matrix of the joint connections $\mathbf{A}\in\mathbb{R}^{J\times J}$ to model the object connectivity. A layer normalisation and skip connection is applied to give the layer output. As with our motion encoder, we stack $O$ layers with residual connections between them.
@@ 3.2.4 Decoder and Procrustes-based loss
Lastly, given the latent features $\mathbf{F}_{S}$, we use an MLP to decode these features to obtain the predicted 3D structures $\mathbf{\widehat{Y}}^{\text{canon}}\in\mathbb{R}^{T\times J\times 3}$ of the object in a canonical 3D space
@@
We align each canonical prediction $\mathbf{\widehat{Y}}^{\text{canon}}$ with the ground truth $\mathbf{Y}$ via a Procrusteas alignment method which solves for the optimal rotation $\mathbf{\tilde{R}}_{t}$ individually for each frame $t$ as:
@@ 3.2.5 Loss function
We use $\lambda$ as a scalar to weight the velocity loss.
@@ 4 Experiments
Table 2: Quantitative comparison between multi-category (MC) and single-category training. We use a ✓for models trained with multi-category training. Each method benefits from multi-category training. See supp. material for a breakdown of per-animal results.
We evaluate our method on various animal categories to assess its performance and generalisation properties. Comparative analyses are with recent state-of-the-art video (MotionBERT) [34] and single-frame (3D-LFM) lifting models [7]. For completeness, we also include comparisons on human-specific lifting, see the supp. material for results on the Human3.6M benchmark.
@@
We use a dataset of 13 animals generated with our data pipeline, as described in Sec. 3. The 2D keypoints provided by off-the-shelf pose detectors [30] are inherently noisy due to factors such as lighting conditions and image quality. To simulate these conditions, we synthetically perturb the 2D keypoints with an additive noise, which corresponds to an average $3$-pixel error. We present results for non-noisy data in the supp. material. Note that we also provide comparisons on a real-world example, see Fig.  and 3D human pose estimation results in the supp. material. We normalise the 2D keypoints and 3D labels to $[-1,1]$, after scaling the 3D labels following existing works on 3D human pose estimation [34]. We split the data for training by randomly selecting 80% of the animations sequences
for each animal, with the remaining animations withheld for testing. For further information about the data, see Sec. 3.
@@
Previous video lifting methods [34, 4, 11, 20] evaluate the non-rigid structure and motion of their method by calculating the mean per-joint position error (MPJPE) directly with the ground truth in the camera space. However, our model and 3D-LFM make predictions in a canonical space that requires the alignment of each frame to the ground truth. As such, we instead use the standard per-frame Procrustes-aligned MPJPE metric and refer to it as the frame-aligned MPJPE (FA-MPJPE) for brevity. Additionally, we compose a metric to measure the relative motion error in a video sequence.
@@
We formulate the sequence-aligned MPJPE (SA-MPJPE) as solving for a single rotation matrix $\mathbf{R}\in SO(3)$ to align the 3D predictions $\mathbf{\widehat{Y}}$ and ground truth $\mathbf{Y}$ for all $T$ frames in a sequence. This is in contrast to the SA-MPJPE that aligns each individual frame in a sequence by solving for $T$ rotation matrices. After alignment we compute the MSE to produce the final error value. Let us specify $\mathbf{\widehat{Y}},\mathbf{Y}\in\mathbb{R}^{T\times J\times 3}$, for $J$ joints, to define our metric as
@@
Compared to Eq. 13 we are instead solving for only one global $\mathbf{R}$. Our metric thus captures any error in the motion that occurs between subsequent frames of a sequence.
Lastly, we are able to report the mean per-joint velocity error (commonly referred to as MPVE) after we align the predictions $\mathbf{\widehat{Y}}$ to the ground truth for the whole sequence. For clarity in our comparisons, we refer to this as SA-MPVE.
@@
Here we provide important implementation details of our method and refer the reader to the supp. material for further details. We train on a total of 871 video sequences and evaluate on a separate set of 199 unseen sequences across all 13 animal categories. We construct batches of 32 sequences with 48 frames per sequence. We zero-pad the inputs up to the maximum of 29 joints that occur in dataset. We choose a layer size of 4 for the motion encoder and 12 for the space encoder. The hidden-dimension size $D$ is 256. Experiments were conducted on a single NVIDIA A100 GPU.
MotionBERT uses human-specific semantic knowledge, making it unsuited for object-agnostic lifting. To provide a fair comparison, we do not modify the proposed architecture and instead apply two other alterations. We first set the amount of learned positional embeddings to the maximum amount of joints seen in the dataset, allowing it to handle all animal rigs. We also randomly permute the 2D inputs during training and testing to simulate a real the object-agnostic scenario. This is required so that the positional embeddings are not learning dataset-specific skeleton semantics. We ensure that joints are permuted the same for a video sequence such that there remains temporal correspondence.
@@ 4.1 Object-agnostic lifting
Tab. 1 demonstrates the effectiveness of our method compared to both MotionBERT and 3D-LFM. We outperform both methods across all three metrics and nearly every animal. Notably, our approach achieves 45% lower SA-MPJPE and 70% lower SA-MPVE compared to 3D-LFM, demonstrating that our predicted 3D motion has both higher accuracy and smoother movement, while also preserving accurate structure (FA-MPJPE). This substantial performance gap highlights the critical role of the motion encoder in capturing the temporal relationship between joints. While MotionBERT is a spatio-temporal approach, it has no inductive bias to assist with handling scarce data and multi-category training. Fig. 2(a) qualitatively demonstrates the predictions of our method compared to 3D-LFM.
@@
We evaluate the performance of each method on single-category training and compare it to their performance on multi-category training in Sec. 4.1. For the single-category training, we train each approach from scratch using data specific to one animal category. This process is repeated for all $13$ animals, and the mean error is reported in Tab. 2. All methods show significant improvement from multi-category training, achieving at least a 25% reduction in FA-MPJPE, with our approach achieving a 50% reduction. This improvement highlights the advantage of unified learning across a vast spectrum of object categories, particularly in scenarios where the training data is small and unbalanced.
@@
To evaluate the robustness of 3D-LFM and our approach to occlusion scenarios, we trained both models by randomly masking $10$% of all 2D keypoints within a frame and tasking the models with recovering the 3D locations of the missing joints. As shown in Fig. 5(a) with $10\%$ occlusion, our method demonstrates superior robustness to occlusions while 3D-LFM struggles. Even at an extreme case of 60% occlusion we continue to see legible predictions from our method, see supp. material for a qualitative result.
@@ 4.2 OOD generalization
We perform a $13$-fold analysis, where each fold involves holding out one animal category from the original dataset during training. For example, in $1$-fold, the bunny category is excluded from the training data and used to evaluate the generalisation capability of each method. As shown in Fig. 3 (left), our approach demonstrates superior OOD generalization when handling unseen animal categories by outperforming existing methods by a significant margin. We present a qualitative reconstruction for a bunny instance in Fig. 4(a) and refer the reader to the supp. material for tabulated results.
@@
When lifting an unseen animal in the wild, we may encounter an animal with a more complex structure than seen during training. We showcase our ability to generalize to an unseen animal with an unseen number of joints. While MotionBERT is limited to rigs with the same or fewer joints as those seen during training, our method can handle any number of joints. We train on animal rigs with $27$ or fewer joints and test on two unseen animals with $29$ joints (deer and moose). As shown in Fig. 3 (right), while both 3D-LFM and our approach are impacted by the difficulty of the task. Our approach outperforms 3D-LFM by at least 40% in all metrics. We show the tabulated results in the supp. material.
@@ 4.3 Ablations
We observe that applying our inductive bias to restrict information sharing enhances the optimization of our model. We conduct our ablation by starting from having a strong inductive bias ($\alpha=2$) inductive bias and progressively increasing it by a factor of $2$. We do this until $\alpha$ is large enough to allow information sharing across the entire sequence ($\alpha=-$). We apply $\alpha$ according to Eq. 7. Tab. 4 shows the impact of different $\alpha$ on performance, averaged over five statistical runs. Overall, we observe a trade-off between the accuracy of the 3D structure and the accuracy of 3D motion. In our case we identify $\alpha=8$ as optimal in our case, which we used for all of our experiments.
@@
Here we ablate our use of our analytical RFF for encoding temporal information. In Tab. 5 we show that in our experiments it is beneficial to use an analytical embedding over a learned embedding. This embedding is important for the motion encoder to understand the ordering of frames in a sequence.
@@
Similar to 3D-LFM, we find it useful to train our model with a Procrustes-based loss (PBL), as shown in Tab. 6. Allowing the model to focus on learning object structures significantly enhances the predicted 3D structure and benefits overall 3D motion accuracy. Interestingly, we observe that we are able to perform well without the PBL, even outperforming 3D-LFM across all metrics. This offers an interesting insight into future work with our method for practical applications that require implicitly predicted camera rotation.
@@
Table 6: Procrustes vs. non-Procrustes-based training. Our spatio-temporal 2D-3D lifting approach with Procrustean alignment surpasses 3D-LFM.
@@ 5 Conclusion
In this work, we introduced an object-agnostic 3D lifting model that leverages a temporal inductive bias for temporal sequences. Our approach sets a new benchmark in lifting performance for object categories with limited available data. The model’s ability to generalize across unseen object categories and rigs shows its versatility and robustness, even in challenging scenarios involving noise and occlusions. In addition to these contributions, we have also introduced a new synthetic dataset designed to stimulate further research in class-agnostic 3D lifting models. This work paves the way for more generalized and efficient 3D reconstruction methods that can be applied across diverse real-world applications.
@@ B.1 Statistics
An overview of our dataset can be seen in Tab. 7. We showcase the number of joints, animations, and frames for each animal category. Our dataset contains is a variety of different categories and rigs to facilitate class-agnostic training and evaluation.

### v2

@@ 3.2.3 Space encoder
The space encoder uses the features from our motion encoder to model the relationships among joints in a single frame. We found the hybrid graph-based approach of [8] to perform favorably for our task. Given the transposed motion features $\mathbf{F}_{M}\in\mathbb{R}^{T\times J\times D}$, a single space-layer has two simultaneous processing streams, one for capturing the local connectivity between joints $G_{\text{local}}$, and another for capturing global connectivity $G_{\text{global}}$. These two streams are concatenated and projected to provide an output of spatial features containing a combination of both streams:
@@
The local graph-attention $G_{\text{local}}$ utilises an adjacency matrix of the joint connections $\mathbf{A}\in\mathbb{R}^{J\times J}$ to model the object connectivity. A layer normalisation and skip connection is applied to produce the final output of a layer. As with our motion encoder, we stack $O$ layers with residual connections between them.
@@ 3.2.4 Decoder and Procrustes-based loss
Lastly, given the latent spatial features $\mathbf{F}_{S}$, we use an MLP to decode the predicted 3D structures of the object in a canonical 3D space
@@
We align each canonical prediction $\mathbf{\widehat{Y}}^{\text{canon}}\in\mathbb{R}^{T\times J\times 3}$ with the ground truth $\mathbf{Y}$ via a Procrustes alignment method which solves for the optimal rotation $\mathbf{\tilde{R}}_{t}$ individually for each frame $t$ as:
@@ 3.2.5 Loss function
We use the scalar $\lambda$ to weight the velocity loss.
@@ 4 Experiments
Table 2: Quantitative comparison between multi-category (MC) and single-category training. We use a ✓for models trained with multi-category training. Each method benefits from multi-category training. See supplementary material for a breakdown of per-animal results.
We evaluate our method on various animal categories to assess its performance and generalisation properties. Comparative analyses are with recent state-of-the-art video (MotionBERT [36]) and single-frame (3D-LFM [8]) lifting models on various animal categories and motion sequences.
@@
We use the AnimalSyn3D dataset as described in Sec. 3.1. The 2D keypoints provided by off-the-shelf pose detectors [32] are inherently noisy due to factors such as lighting conditions and image quality. To simulate these conditions, we synthetically perturb the 2D keypoints with an additive Gaussian noise, which corresponds to a $3$-pixel error on average. We present results for non-noisy data (Tab. 8) and also provide comparisons on 3D human pose estimation in the supplementary material (Tab. 11), although human-specific lifting is not the focus of our work. We normalise 2D keypoints and 3D labels to $[-1,1]$, after scaling the 3D labels following existing works on 3D human pose estimation [36]. We split the data for training by randomly selecting 80% of the animation sequences for each animal, with the remaining animations withheld for testing.
@@
Previous video lifting methods [36, 5, 12, 21] evaluate the non-rigid structure and motion of their approach by calculating the mean per-joint position error (MPJPE) directly with the ground truth in the camera space. However, our model and 3D-LFM make predictions in a canonical space that requires the alignment of each frame to the ground truth. As such, we instead use the standard per-frame Procrustes-aligned MPJPE metric and refer to it as the frame-aligned MPJPE (FA-MPJPE) for brevity. Additionally, we compose a metric to measure the relative motion error in a video sequence.
@@
We formulate the sequence-aligned MPJPE (SA-MPJPE) as solving for a single rotation matrix $\mathbf{R}\in SO(3)$ to align the 3D predictions $\mathbf{\widehat{Y}}$ and ground truth $\mathbf{Y}$ for all $T$ frames in a sequence. This is in contrast to the SA-MPJPE that aligns each individual frame in a sequence by solving for $T$ rotation matrices. After alignment, we compute the MSE to produce the final error value. Let us specify $\mathbf{\widehat{Y}},\mathbf{Y}\in\mathbb{R}^{T\times J\times 3}$, for $J$ joints, to define our metric as
@@
Compared to Eq. 13, we are instead solving for a single, global $\mathbf{R}$. Our metric thus captures any error in the motion that occurs between subsequent frames of a sequence.
Lastly, we are able to report the commonly used mean per-joint velocity error (MPVE) after performing the global sequence alignment. For clarity in our comparisons, we refer to this as SA-MPVE.
@@
Here we provide important implementation details of our method and refer the reader to the supplementary material for further details. We construct batches of 32 sequences with 48 frames per sequence. We train on a total of 871 video sequences and evaluate on a separate set of 199 unseen sequences across all 13 animal categories. Inputs are zero-padded up to the maximum of 29 joints that occur in dataset. A layer size of $P=4$ is chosen for the motion encoder and $O=12$ for the space encoder. The hidden-dimension size $D$ is 256. Experiments were conducted on a single NVIDIA A100 GPU.
MotionBERT uses human-specific semantic knowledge, making it unsuited for object-agnostic lifting. To provide a fair comparison, we do not modify the proposed architecture and instead apply two alternative alterations. We first set the number of learned positional embeddings to the maximum amount of joints seen in the dataset, allowing it to handle all animal rigs. We also randomly permute the 2D inputs during training and testing to simulate a real object-agnostic scenario. This is required so that the positional embeddings are not learning dataset-specific skeleton semantics. We ensure that joints being permuted retain their temporal correspondence over a sequence.
@@ 4.1 Object-agnostic lifting
Tab. 1 demonstrates the effectiveness of our method compared to MotionBERT and 3D-LFM. We outperform both methods across all three metrics and nearly every animal. Notably, our approach achieves 45% lower SA-MPJPE and 70% lower SA-MPVE compared to 3D-LFM, demonstrating that our predicted 3D motion has more accurate and smoother movement, while also preserving high-fidelity object structure (FA-MPJPE). This substantial performance gap highlights the critical role of the motion encoder in capturing the temporal relationship of joints. While MotionBERT is a spatio-temporal approach, it has no inductive bias to assist with handling scarce data and multi-category training. Fig. 3(a) qualitatively demonstrates the predictions of our method compared to 3D-LFM.
@@
We evaluate performance on single-category training and compare it to their performance on multi-category training. For single-category, we train each approach from scratch using data specific to a single animal category. This process is repeated for all $13$ animals, and the mean error is reported in Tab. 2. All methods show significant improvement when performing multi-category training as opposed to single-category, achieving at least a 25% reduction in FA-MPJPE, with our approach achieving a 50% reduction. This improvement highlights the advantage of unified learning across a vast spectrum of object categories, particularly in scenarios where the training data is small and unbalanced.
@@
To evaluate the robustness of 3D-LFM and our approach to occlusion scenarios, we trained both models by randomly masking $10$% of all 2D keypoints within a frame and tasking the models with recovering the 3D locations of the missing joints. As shown in Fig. 6, our method demonstrates superior robustness while 3D-LFM struggles. Even in an extreme case of 60% occlusion we continue to see legible predictions from our method; see Fig. 8 in the supplementary material for a qualitative result.
@@ 4.2 OOD generalization
We perform a $13$-fold analysis, where each fold involves holding out one animal category from the original dataset during training. For example, the bunny category is excluded from the training data and used to evaluate the generalization capability of each method. As shown in Fig. 4 (left), our approach demonstrates superior OOD generalization when handling unseen animal categories by outperforming existing methods by a significant margin. We present a qualitative reconstruction for a bunny instance in Fig. 5(a) and refer the reader to Tab. 9 in the supplementary material for tabulated results.
@@
When lifting an unseen animal in the wild, we may encounter an animal with a more complex structure than seen during training. We showcase our ability to generalize to an unseen animal with an unseen number of joints. While MotionBERT is limited to rigs with the same or fewer joints as those seen during training, our method can handle any number of joints. We train on animal rigs with $27$ or fewer joints and test on two unseen animals with $29$ joints (deer and moose). As shown in Fig. 4 (right), while both 3D-LFM and our approach are impacted by the difficulty of the task, our approach outperforms 3D-LFM by at least 40% in all metrics. We show the tabulated results in the supplementary material (Tab. 10).
@@ 4.3 Ablations
We observe that applying our inductive bias to restrict information sharing enhances the optimization of our model. We conduct our ablation by progressively increasing the window size $\alpha$. We apply $\alpha$ according to Eq. 7. Tab. 4 shows the impact of different $\alpha$ on performance, averaged over five statistical runs. Overall, we observe a trade-off between the accuracy of the 3D structure and the accuracy of 3D motion. In our case, we identify $\alpha=8$ as optimal, which we used for all of our other experiments.
@@
Here we ablate our use of analytical RFF to encode temporal information. In Tab. 5, we show that it is beneficial to use an analytical embedding over a learned embedding. Analytical RFF provides a significant performance increase over a learned embedding, in our case it seems to result from a scarcity of data.
@@
Similar to 3D-LFM, we find it useful to train our model with a Procrustes-based loss, as shown in Tab. 6. Allowing the model to focus solely on learning object structure significantly enhances the predicted 3D structure and benefits overall 3D motion accuracy. Interestingly, we observe that we are able to perform well without the Procrustes-based loss, even outperforming 3D-LFM across all metrics. This offers an interesting insight into future work involving our method for practical applications that require implicitly predicted camera rotation.
@@
Table 6: Procrustes vs. non-Procrustes training. Our spatio-temporal 2D-3D lifting approach with Procrustean alignment surpasses 3D-LFM.
@@ 5 Conclusion
In this work, we introduced an object-agnostic 3D lifting model that leverages a temporal inductive bias for temporal sequences. Our approach sets a new benchmark in lifting performance for object categories with limited available data. The model’s ability to generalize across unseen object categories and rigs shows its versatility and robustness, even in challenging scenarios involving noise and occlusions. In addition to these contributions, we have introduced a new synthetic dataset, AnimalSyn3D, designed to stimulate further research in class-agnostic 3D lifting models. This work aims to pave the way for more generalized and efficient 3D reconstruction methods that can be applied across diverse real-world applications.
6 Acknowledgements
We would like to thank Angel Allen for contributing the dog and raccoon to the dataset, Hemanth Saratchandran for polishing the mathematical notations, Lachlan Mares and Stefan Podgorski for notes during drafting, and Chee-Keng Ch’ng for helpful discussions throughout the project.
@@ B.1 Statistics
An overview of our dataset can be seen in Tab. 7. We showcase the number of joints, animations, and frames for each animal category. Our dataset contains a variety of different categories and rigs to facilitate class-agnostic training and evaluation.